OpenStudy (anonymous):

Need help expanding x^3/(x^2+1)^2 into partial fractions.

OpenStudy (anonymous):

\[\frac{x^3}{(x^2+1)^2}=\frac{Ax + B}{x^2 +1} + \frac{Cx+D}{(x^2+1)^2}\]

OpenStudy (anonymous):

\[\huge\frac{x^3}{(x^2+1)^2}=\frac{Ax + B}{x^2 +1} + \frac{Cx+D}{(x^2+1)^2}\]

OpenStudy (anonymous):

Isn't it of the form \[(ax+b)^k\]?

OpenStudy (anonymous):

I'm not sure what you mean. The original or what i've done?

OpenStudy (anonymous):

I mean why is it expanded as \[(ax^2+bx+c)^k\]

OpenStudy (anonymous):

because it matches that form if you allow the b coefficient to equal zero

OpenStudy (anonymous):

Okay. Thanks a lot.

OpenStudy (anonymous):

The order of the numerator will always lag behind the highest order of factor